OpenStudy (anonymous):

I just want to see if some can check my work to make sure I'm completing these problems right. The problem states evaluate the definite integral of each function.

OpenStudy (anonymous):

\[\int\limits_{1}^{7}(6x^2+2x-3)dx\] \[\int\limits_{1}^{7}6x^2dx +\int\limits_{1}^{7}2xdx -\int\limits_{1}^{7}3dx\] \[6\int\limits_{1}^{7}x^2dx=2x^3+ 2\int\limits_{1}^{7}xdx -\int\limits_{1}^{7}3dx\] \[=2x^3+x^2-3x+c\] \[=x(2x^2+x-3)+c\] \[=7(2(7)^2+7-3)=714\]

OpenStudy (zzr0ck3r):

the answer is correct but you have things that dont make sense \(6\int\limits_{1}^{7}x^2dx=2x^3+ 2\int\limits_{1}^{7}xdx -\int\limits_{1}^{7}3dx\)

OpenStudy (zzr0ck3r):

is one example^^^

OpenStudy (midhun.madhu1987):